OpenStudy (kellyspeakslouder):

f(x) = Square root of quantity x plus seven.; g(x) = 8x - 11 Find f(g(x)). A. f(g(x)) = 2Square root of quantity two x plus one. B. f(g(x)) = 8Square root of quantity x plus seven. - 11 C. f(g(x)) = 8Square root of quantity x minus four. D. f(g(x)) = 2Square root of quantity two x minus one.

OpenStudy (kellyspeakslouder):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

This is f(x) right? \[\Large f(x) = \sqrt{x+7}\] The 7 is under the square root?

OpenStudy (kellyspeakslouder):

correct :)

jimthompson5910 (jim_thompson5910):

Replace every copy of 'x' with g(x). Then replace the g(x) on the right side with 8x-11 \[\Large f(x) = \sqrt{x+7}\] \[\Large f({\color{red}{x}}) = \sqrt{{\color{red}{x}}+7}\] \[\Large f({\color{red}{g(x)}}) = \sqrt{{\color{red}{g(x)}}+7}\] \[\Large f({\color{red}{g(x)}}) = \sqrt{{\color{red}{8x-11}}+7}\] I'll let you simplify

OpenStudy (kellyspeakslouder):

would the square root be.. 8x minus 4? so C?

jimthompson5910 (jim_thompson5910):

what does 8x-11+7 simplify to?

OpenStudy (kellyspeakslouder):

I keep getting 8x-4. I did -11+7 and got -4

jimthompson5910 (jim_thompson5910):

\[\Large f(g(x)) = \sqrt{8x-11+7}\] \[\Large f(g(x)) = \sqrt{8x-4}\] \[\Large f(g(x)) = \sqrt{4(2x-1)}\] \[\Large f(g(x)) = \sqrt{4}*\sqrt{2x-1}\] I'll let you do the last step

OpenStudy (kellyspeakslouder):

-8?

jimthompson5910 (jim_thompson5910):

what is the square root of 4?

OpenStudy (kellyspeakslouder):

2

jimthompson5910 (jim_thompson5910):

so \[\Large f(g(x)) = \sqrt{4}*\sqrt{2x-1}\] \[\Large f(g(x)) = 2\sqrt{2x-1}\]

OpenStudy (kellyspeakslouder):

so D?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (kellyspeakslouder):

awesome sauce! thanks!

jimthompson5910 (jim_thompson5910):

sure thing